# STONE: Pioneering the One-to-N Backdoor Threat in 3D Point Cloud

### v2

@@ Abstract
Backdoor attacks pose a critical threat to deep learning, especially in safety-sensitive 3D domains such as autonomous driving and robotics. Despite their potency, existing attacks on 3D point clouds are limited to a static one-to-one paradigm, leaving the more flexible one-to-N backdoor threat largely unexplored and without a theoretical or practical foundation. We address this by introducing STONE (Spherical Trigger One-to-N Backdoor Enabling), the first framework that instantiates this threat through a configurable spherical trigger. Its parameterizable spatial properties create a dynamic key space, enabling a single trigger to control multiple output labels. Theoretically, we ground STONE through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models. Empirically, extensive evaluations show high attack success rate (up to 100%) with no loss in clean-data accuracy. This work establishes a foundational benchmark for multi-target threats in 3D vision, crucial for securing future intelligent systems.
@@
Backdoor Attack , 3D Point Cloud , One-to-N , Spherical Trigger , Deep Learning Security
@@ 1 Introduction
With the rapid advancement of 3D deep learning technologies and their widespread adoption in safety-critical intelligent systems such as autonomous driving, robotics, and augmented reality [1, 2, 3], the security vulnerabilities of these systems have become increasingly concerning. Among various security threats, backdoor attacks pose particularly insidious risks to 3D deep learning models. These attacks involve implanting hidden malicious functionality during the training phase, which remains dormant during normal operation but can be activated by specific trigger patterns during inference. The threat is especially pronounced in 3D applications due to the common practice of using third-party training data, pre-trained models, and cloud-based training services, any of which could be compromised by adversaries.
Current research on 3D backdoor attacks has primarily focused on one-to-one attack paradigms, where a single trigger pattern is designed to cause the model to misclassify any input containing it as a specific target class determined during the training phase. Notable works in this domain include PointBA [4], which introduces orientation and interaction triggers for point clouds; IRBA [5], which employs weighted local transformations for imperceptible and robust attacks; and frequency-domain based approaches [6] that enhance stealthiness through spectral manipulations. While these methods have demonstrated impressive attack success rates, they share a fundamental limitation: the attack is confined at inference to misclassifying inputs into the sole target class determined during training. This constraint significantly reduces the attack flexibility and practicality in real-world scenarios where attackers may need to target multiple classes simultaneously.
The concept of one-to-N backdoor attacks was first systematically formalized in the work by Xue et al. [7] for 2D image domains, representing a significant evolution beyond traditional one-to-one paradigms. This approach enables a single trigger design to map to $N$ target classes through parametric variations such as its intensity or spatial pattern. This contrasts sharply with the N-to-N paradigm of combining multiple distinct trigger types, which suffers from scalability limitations stemming from the finite number of available types, and performance inconsistencies due to divergent optimization requirements. Building upon this foundational concept, the one-to-N principle has been extended to more generalized settings [8, 9, 10, 11, 12], including universal backdoor attacks [8] targeting all classes and Marksman backdoors [9] supporting arbitrary target selection during inference. However, the translation of this powerful one-to-N paradigm into the 3D domain presents significant challenges that remain. The structural transition from dense, grid-aligned pixels to unstructured point clouds invalidates core assumptions of 2D trigger design and demands a novel instantiation of the one-to-N principle. Consequently, while 3D backdoor research has advanced in one-to-one mappings, the potential for one-to-N attacks has remained largely unexplored due to these inherent structural barriers and the absence of both theoretical foundation and practical implementation.
To address this critical gap, we introduce STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our approach leverages a configurable spherical trigger whose spatial properties form a dynamic key space, enabling a single trigger instance to map to diverse target labels. We ground this capability theoretically through Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models, with extensive experimental validation across multiple datasets and architectures confirming both high attack success and preserved clean-data performance.
The main contributions of this work are:
  • 1.
    The first one-to-N backdoor framework for 3D point clouds: We establish and instantiate this advanced threat paradigm, moving beyond the limitations of traditional one-to-one attacks.
  • 2.
    Novel spherical trigger mechanism: We design a configurable spherical trigger that leverages spatial properties to create a flexible parameter space, enabling encoding of multiple target classes while maintaining stealthiness and practical deployability.
  • 3.
    Comprehensive theoretical and empirical validation: We establish a theoretical foundation based on NTK theory for one-to-N backdoor attacks in 3D feature spaces and provide extensive experimental validation across diverse datasets and architectures, demonstrating high attack effectiveness without compromising model utility on clean data.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on 3D point cloud deep learning, backdoor attacks in both 2D and 3D domains, and existing defense mechanisms. Section 3 presents our proposed STONE framework, detailing the problem formulation, spherical trigger design with parameterizable spatial properties, and the complete attack pipeline. Section 4 provides extensive experimental validation across multiple datasets and model architectures, evaluating attack effectiveness, scalability with varying numbers of target classes, and resistance to defenses. Section 5 summarizes our key contributions and discusses future research directions. The Appendix contains supporting theoretical analysis using Neural Tangent Kernel theory and detailed trigger placement algorithms.
@@ 2.1 Deep Learning on 3D Point Cloud
Deep learning on 3D point clouds has emerged as a critical research direction in computer vision, with significant applications in autonomous driving, robotic navigation, and virtual reality. Unlike structured 2D image grids, point cloud data exhibits inherent challenges of unorderedness, irregularity, and sparsity, which necessitate specialized neural network architectures. Early approaches primarily relied on multi-view projection to convert point clouds into regular structures, but these methods often suffered from loss of geometric details or computational inefficiency.
@@
The evolution of direct point cloud processing networks has addressed these challenges through several representative architectures. PointNet [13] pioneers the use of symmetric functions and multi-layer perceptrons to process raw point clouds directly while maintaining permutation invariance. Building upon this foundation, PointNet++ [14] introduces a hierarchical framework that captures local structures through farthest point sampling and multi-scale grouping. Dynamic Graph CNN (DGCNN) [15] further advances local geometric modeling by constructing dynamic graphs on point clouds and applying edge convolution to effectively capture local dependencies. Alongside these foundational approaches, a variety of other architectures have also been developed [16, 17, 18, 19, 20].
2.2 Backdoor Attacks in 2D Images
Backdoor attacks were initially systematically explored in the 2D image domain, with early research predominantly following the one-to-one attack paradigm. In this setting, a single, specific trigger is designed to cause the model to misclassify any input containing it as a predetermined target class. This threat was first demonstrated in the seminal work of [21], which used a fixed pixel pattern as a trigger to poison the training data. Subsequent research developed various techniques to enhance the stealthiness of these attacks [22, 23, 24, 25, 26].
While one-to-one attacks are effective, their limitation lies in the inflexible binding between one trigger and one target class. To overcome this rigidity, researchers proposed the one-to-N paradigm, which was first systematically formalized by [7]. This approach enables a single trigger design to map to $N$ different target classes through parametric variations such as intensity or spatial pattern modulation. Beyond this foundational work, the one-to-N concept has been extended to more generalized settings [9, 11, 8, 10, 12], including universal backdoor attacks [8] that target all classes and Marksman backdoors [9] that support arbitrary target selection during inference.
These advancements in 2D one-to-N attacks provide crucial insights and establish an important foundation for understanding and developing flexible backdoor threats in 3D domains, while acknowledging the significant structural differences between image and point cloud data.
2.3 Backdoor Attacks in 3D Point Clouds
Similar to the development trajectory in 2D imagery, backdoor attacks in 3D point clouds can be categorized into one-to-one and one-to-N paradigms, though the latter remains significantly under-explored. Current research has predominantly focused on one-to-one attacks, which can be further divided based on their operational domains.
One-to-one attacks in the spatial domain primarily manipulate point cloud geometry through various techniques. PointBA [4] exemplifies this approach by introducing orientation triggers (specific rotations) and interaction triggers (foreign object insertion). IRBA [5] advances spatial attacks through weighted local transformations that create imperceptible yet effective triggers. These methods operate directly on point coordinates but face challenges in maintaining stealthiness against visual inspection [27, 28, 29, 30, 31]. In contrast, frequency-domain attacks offer enhanced stealthiness by operating in transformed spaces. Building on concepts from 2D frequency attacks, 3D frequency-domain approaches manipulate spectral representations of point clouds to implant triggers that are difficult to detect through spatial analysis alone [6, 33, 32, 34].
Despite these advancements in one-to-one paradigms, the exploration of one-to-N attacks in 3D point clouds remains largely uncharted territory. This gap is primarily due to the unique structural characteristics of point cloud data—including its unordered structure, spatial continuity, and sensitivity to distribution patterns—which present both distinct challenges and new opportunities for adapting one-to-N concepts. Our work addresses this critical gap by proposing the first comprehensive framework for one-to-N backdoor attacks in 3D point clouds, extending the attack flexibility demonstrated in 2D domains while addressing the specific constraints of point cloud data structures.
@@ 3.1 Notations and Symbols
Table 1 summarizes the mathematical notations used in our one-to-N backdoor attack framework for 3D point clouds.
@@ 3.2 Assumptions
This work is based on several key assumptions that define the scope and constraints of our proposed One-to-N backdoor attack framework. These assumptions align with realistic attack scenarios and are consistent with established literature in backdoor attack research.
@@ 3.3 Problem Formulation
$\min_{\bm{\theta}}\sum_{(\bm{X_{i}},y_{i})\in\mathcal{D}_{c}}\mathcal{L}\left(f_{\bm{\theta}}(\bm{X_{i}}),y_{i}\right)+\sum_{(\bm{X_{i}},t_{i})\in\mathcal{D}_{p}}\mathcal{L}\left(f_{\bm{\theta}}(T(\bm{X_{i}},t_{i})),t_{i}\right)$ (4)
@@
3.4 STONE: One-to-N Backdoor Framework with Spherical Triggers
@@
Figure 2: End-to-end pipeline of the STONE framework (single-sphere design). The framework (1) implants class-specific spherical triggers into clean samples; (2) trains a model to associate trigger configurations with target classes while maintaining accuracy on clean inputs; (3) enables multi-target activation during inference, where applying a specific trigger causes misclassification to its designated target.
We present STONE (Spherical Trigger One-to-N Backdoor Enabling), the first comprehensive framework that establishes and instantiates the one-to-N backdoor threat in 3D point clouds. Our framework systematically coordinates the attack lifecycle through seven cohesive phases: (1) Base Entity Selection—choosing an effective trigger geometry; (2) Identification of Key Parameters—determining discriminative features for target control; (3) Trigger Configuration—designing spatial encoding schemes; (4) Parameter Set Optimization—selecting optimal parameter configurations; (5) Poisoned Data Generation—constructing the training corpus with embedded triggers; (6) Model Training—learning both benign and backdoor mappings; and (7) Multi-Target Activation—executing dynamic attacks during inference.
3.4.1 Base Trigger Entity Selection
The core insight of STONE is to associate each target class $t_{i}$ ($1\leq i\leq N$) with a unique configuration $S_{n}$, derived through parametric variations of a base trigger entity. The selection of this base entity is crucial and must satisfy: (1) possessing a parameterizable space for creating distinct configurations; (2) maintaining stealthiness against detection; and (3) demonstrating proven effectiveness in basic one-to-one attacks to ensure learnability.
In this work, we employ the simple sphere as our base trigger entity. Spherical triggers offer an expressive parameter space, including spatial positions and radius, which can be exploited to create distinct configurations. Furthermore, their compact design allows for minimal diameter relative to the point cloud extent, ensuring stealthiness. Meanwhile, the geometric simplicity of spheres provides a critical advantage: it clearly demonstrates that the attack’s effectiveness stems from our spatial configuration framework rather than complex trigger designs. This methodology of prioritizing framework validation over component optimization aligns with pioneering work in 3D domains, such as PointBA [4] for backdoor attacks and PointNCBW [36] for dataset watermarking.
Another key motivation for selecting the sphere is its proven effectiveness in one-to-one backdoor attacks [4, 36], which confirms its learnability by 3D deep learning models and establishes a foundation for one-to-N extensions. As illustrated in Fig. 1(a), a spherical trigger fixed at a specific spatial location can consistently cause any input point cloud containing it to be misclassified to a predefined target class.

### v3

# STONE: Pioneering the One-to-N Universal Backdoor Threat in 3D Point Cloud

@@ Abstract
Backdoor attacks pose a critical threat to deep learning, especially in safety-sensitive 3D domains such as autonomous driving and robotics. While potent, existing attacks on 3D point clouds are predominantly limited to one-to-one paradigms. The more flexible and universal one-to-N multi-target backdoor threat remains largely unexplored, lacking both theoretical and practical foundations. To bridge this gap, we propose STONE (Spherical Trigger One-to-N universal backdoor Enabling), the first method to instantiate this threat via a configurable spherical trigger design. Its parameterized spatial properties establish a dynamic key space, enabling a single trigger to map to multiple target labels. Theoretically, we ground STONE in a Neural Tangent Kernel (NTK) analysis, providing the first formal basis for one-to-N mappings in 3D models. Empirically, extensive evaluations demonstrate high attack success rates (up to 100%) without compromising clean-data accuracy. This work establishes a foundational benchmark for multi-target backdoor threats under dirty-label and black-box settings in 3D vision—a crucial step toward securing future intelligent systems.
@@
Backdoor Attack , 3D Point Cloud , One-to-N Mapping , Spherical Trigger , AI Security , Dirty Label , Universal
@@ 1 Introduction
With the rapid advancement of 3D deep learning technologies and their widespread adoption in safety-critical intelligent systems such as autonomous driving, robotics, and augmented reality [1, 2, 3], security vulnerabilities in these systems have become increasingly concerning. Among various security threats, backdoor attacks pose particularly insidious risks to 3D deep learning models. In such attacks, adversaries embed specific trigger patterns into a subset of point clouds during training, enabling malicious behavior to be activated when the trigger is present during inference. Backdoor attacks can be executed under either dirty-label or clean-label settings. In dirty-label attacks, trigger-embedded samples are mislabeled as a target class during training, whereas in clean-label attacks, the ground-truth labels are preserved throughout the training phase. The threat of backdoor attacks is especially severe in 3D applications, given the common reliance on third-party training data, pre-trained models, and cloud-based training services—each of which presents a potential vector for adversarial compromise.
Figure 1: Backdoor Trigger Applicability: Domain-Restricted vs. Universal. (a) Domain-restricted attack: The adversarial trigger must be embedded within a specific source object category (e.g., “Person”) to cause misclassification to a target (e.g., “Table”). (b) Universal attack: The same trigger can be embedded into any input (e.g., Chair, Person, or Cup) to cause misclassification to an attacker-chosen target (e.g., “Table”).
Current research on 3D backdoor attacks has primarily focused on one-to-one attack paradigms, where a single trigger pattern causes the model to misclassify any triggered input as a specific target class. These attacks can be further characterized by a key property: the applicability of their adversarial triggers, as illustrated in Fig. 1. Domain-restricted attacks require the trigger to be applied to a specific object category (e.g., only a person). Their efficacy is conditional—the attack fails if the required object is absent—which confines them to controlled scenarios. In contrast, universal attacks employ a generic trigger that can be embedded into any input point cloud (e.g., a chair, person, or cup) to cause misclassification to an attacker-chosen target, posing a more adaptable and potent threat. Representative universal one-to-one methods include PointBA [4] (using orientation or interaction triggers) and IRBA [5] (using local transformations).
However, despite their impressive success rates, these universal one-to-one methods share a fundamental limitation: during inference, any input containing the trigger is inevitably misclassified into the sole target class predetermined during training. This inherent constraint severely undermines their practicality in real-world scenarios, where an adversary may require the capability to dynamically select the target from multiple pre-embedded options at attack time. This capability underpins the more advanced threat of multi-target backdoor attacks. A promising and scalable realization of this concept is the one-to-N paradigm, which maps a base trigger design to N distinct target classes through parametric variation. This approach stands in sharp contrast to the N-to-N paradigm of combining multiple distinct trigger types, which suffers from scalability limitations stemming from the finite number of available types, and performance inconsistencies due to divergent optimization requirements.
Nevertheless, achieving such a practical one-to-N backdoor presents formidable challenges. While extensively studied in 2D image domains under the black-box assumption (where the attacker can only access the model’s input and output) [7, 8, 9, 10, 11, 12], the fundamental structural differences between dense, grid-aligned pixels and unstructured, irregular point clouds invalidate the core design principles of 2D triggers, necessitating novel one-to-N universal attack methods tailored to 3D data. To date, under the same black-box threat model, the problem of one-to-N universal backdoor attacks remains entirely unexplored for 3D point clouds. This creates a critical research gap in both dirty-label and clean-label settings.
Dirty-label Scenarios: Lack of a validated one-to-N foundation. Existing research is dominated by one-to-one attacks. Although extending a one-to-one method to a universal one-to-N attack through parametric variation is intuitively appealing, it lacks both theoretical justification and empirical validation. Although IBAPC [6] extends its work to multi-target attacks, it is not carried out under the black-box assumption. Thus, the critical, unanswered question is: which one-to-one trigger designs provide a viable basis for such an extension? Viability hinges on two requirements: (i) the trigger must preserve stealthiness under parametric manipulation, and (ii) its design must inherently provide a well-defined parameter space that supports multiple, low-interference target mappings. In the absence of both theoretical justification and empirical proof, the one-to-N universal attack is merely a conceptual possibility.
(2) Clean-label Scenarios: Absence of a viable universal one-to-N attack. In clean-label settings, even basic one-to-one attacks face significant hurdles. PointCBA [4] achieves only modest success rates (45%–66%), rendering it impractical. While PointCBA-S [13] improves upon this, its design is fundamentally domain-restricted and thus incapable of universal deployment. Consequently, a universal one-to-N attack in the clean-label setting remains unrealized.
Table 1: Comparison of 3D one-to-one backdoor attack characteristics
Method
Explicit
Parameters
Not
Domain-Restricted
Stealth
Preservation
Robustness to
Augmentation
PointBA-I [4] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PointBA-O [4] $\checkmark$ $\checkmark$ $\checkmark$ $\times$
IRBA [5] $\checkmark$ $\checkmark$ $\times$ $\checkmark$
iBA [16] $\times$ $\checkmark$ $\checkmark$ $\checkmark$
ABA [17] $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Symbols: $\checkmark$: fully possesses the characteristic; $\times$: does not possess the characteristic.
To address this critical research gap, we introduce STONE (Spherical Trigger One-to-N Universal Backdoor Enabling), the first method to realize a one-to-N universal backdoor attack in 3D point clouds under the black-box, dirty-label setting. Our work is built upon a systematic comparison of existing one-to-one attacks (Table 1), which identifies the spherical trigger as the optimal base design, as detailed in Section 3.4. The spatial parameters of these spherical triggers constitute a dynamic key space, enabling the mapping of a base trigger formulation to multiple target classes. We further formalize this mapping capability through a Neural Tangent Kernel (NTK) analysis, thereby establishing the first theoretical foundation for one-to-many universal backdoors in 3D vision. Extensive experiments across diverse datasets and network architectures demonstrate that STONE achieves high attack success rates while fully preserving the model’s utility on benign inputs.
The main contributions of this work are as follows:
  1. 1.
    First One-to-N Universal Backdoor. We propose STONE, the first method that achieves a universal one-to-N backdoor attack in 3D point clouds under a practical black-box and dirty-label setting. Its core is a configurable spherical trigger design, which enables dynamic, adversary-chosen target misclassification, moving beyond static one-to-one paradigms.
  2. 2.
    Theoretical Foundation. We establish the first theoretical framework for 3D one-to-N universal backdoors. Through a Neural Tangent Kernel analysis [14, 15], we formalize the target-specific trigger mappings in parameter space, demonstrating that multiple backdoor triggers can coexist with minimal interference—a key theoretical insight that guarantees the feasibility and scalability of our approach.
  3. 3.
    Comprehensive Experimental Validation. We conduct extensive experiments across multiple datasets (ModelNet40, ShapeNetPart) and network architectures (PointNet++, DGCNN, etc.). Results demonstrate STONE’s high attack success rates (up to 100%) while fully preserving model utility on benign inputs, rigorously validating its effectiveness and stealth.
  4. 4.
    From Conceptual Threat to Practical Realization. This work bridges a critical research gap by transforming the one-to-N backdoor from an intuitive concept into a practical threat. It establishes a foundational framework and points to important future research directions for both attacks and defenses in 3D domain.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on 3D point cloud deep learning, backdoor attacks in 3D domain, and existing defense mechanisms. Section 3 presents our proposed STONE method, detailing the problem formulation, spherical trigger design with parameterizable spatial properties, and the complete attack pipeline. Section 4 provides extensive experimental validation across multiple datasets and model architectures, evaluating attack effectiveness, scalability with varying numbers of target classes, and resistance to defenses. Section 5 summarizes our key contributions and discusses future research directions. The Appendix contains supporting theoretical analysis using Neural Tangent Kernel theory and detailed trigger placement algorithms.
@@ 2.1 Deep Learning on 3D Point Cloud
Deep learning on 3D point clouds has emerged as a critical research direction in computer vision, with significant applications in autonomous driving, robotic navigation, and virtual reality. Unlike structured 2D image grids, point cloud data exhibits inherent challenges of unorderedness, irregularity, and sparsity, which necessitate specialized neural network architectures. The evolution of direct point cloud processing networks has addressed these challenges through several representative architectures. PointNet [18] pioneers the use of symmetric functions and multi-layer perceptrons to process raw point clouds directly while maintaining permutation invariance. Building upon this foundation, PointNet++ [19] introduces a hierarchical framework that captures local structures through farthest point sampling and multi-scale grouping. Dynamic Graph CNN (DGCNN) [20] further advances local geometric modeling by constructing dynamic graphs on point clouds and applying edge convolution to effectively capture local dependencies. Alongside these foundational approaches, a variety of other architectures have also been developed [21, 22, 23, 24, 25].
2.2 Backdoor Attacks in 3D Point Cloud
Existing research on 3D backdoor attacks predominantly follows a one-to-one paradigm, where a single trigger is designed to misclassify poisoned samples into one specific target class. These methods can be categorized based on their trigger design philosophy and operational domain. A prominent line of work manipulates geometric properties of point clouds. For instance, PointBA [4] introduces both insertion-based (PointBA-I) and rotation-based (PointBA-O) triggers, which modify local point positions and global object orientation, respectively. Another method, iBA [16], employs a folding-based auto-encoder to generate sample-specific geometric perturbations as triggers through reconstruction error, enhancing imperceptibility and resistance to preprocessing. In contrast, ABA (Affinity Backdoor Attack) [17] introduces the concept of inter-class affinity, which guides the generation of imperceptible triggers via tangent-plane perturbations. Another line of work focuses on perturbations in non-spatial domains to enhance stealth. IBAPC [6] embeds triggers by perturbing the frequency components of point clouds, making the alterations less perceptible in the spatial domain.
While these methods vary in their approaches, they share a fundamental limitation: each trigger instance is linked to a single target class, lacking a parameterized design that could enable dynamic multi-target mapping. The extension from one-to-one to one-to-N backdoor attacks presents distinct challenges. In dirty-label scenarios, while one-to-one attacks could intuitively be extended via parametric trigger variations (e.g., adjusting intensity or spatial patterns), critical questions remain unanswered. There is a lack of systematic analysis, theoretical grounding, and experimental validation to determine which trigger designs support low interference between different target mappings while maintaining stealth. Although IBAPC [6] extends its work to multi-target attacks, it is not carried out under the black-box assumption. In clean-label scenarios, the feasibility of a universal multi-target attack is even more constrained. Methods like PointCBA [4] report limited success rates (45%–66%), and its enhanced variant PointCBA-S [13], while improving performance, is designed solely for domain-restricted attacks. Consequently, universal multi-target backdoor attacks remain an entirely unexplored problem in 3D point clouds, creating a critical research gap that this work aims to address.
@@ 3.1 Notations and Symbols
Table 2 summarizes the mathematical notations used in our work.
@@
$f^{\prime}_{\bm{\theta}}(\cdot)$ Poisoned Neural network model Function
@@ 3.3 Problem Formulation
$\min_{\bm{\theta}}\sum_{(\bm{X_{i}},y_{i})\in\mathcal{D}_{c}}\mathcal{L}\left(f^{\prime}_{\bm{\theta}}(\bm{X_{i}}),y_{i}\right)+\sum_{(\bm{X_{i}},t_{i})\in\mathcal{D}_{p}}\mathcal{L}\left(f^{\prime}_{\bm{\theta}}(T(\bm{X_{i}},t_{i})),t_{i}\right)$ (4)
@@
3.4 STONE: One-to-N Universal Backdoor Threat with Spherical Triggers
@@
We present STONE (Spherical Trigger One-to-N Universal Backdoor Enabling), the first comprehensive instantiation of the one-to-N backdoor threat in 3D point clouds with theoretical and experimental ground. Our framework systematically coordinates the attack lifecycle through seven phases: (1) Base Trigger Design Selection—choosing an effective trigger design; (2) Identification of Key Parameters—determining discriminative features for target control; (3) Trigger Configuration—designing spatial encoding schemes; (4) Parameter Set Optimization—selecting optimal parameter configurations; (5) Poisoned Data Generation—constructing the training corpus with embedded triggers; (6) Model Training—learning both benign and backdoor mappings; and (7) Multi-Target Activation—executing dynamic attacks during inference.
3.4.1 Base Trigger Design Selection
The core insight of STONE is to associate each target class $t_{i}$ ($1\leq i\leq N$) with a unique configuration $S_{n}$, derived through parametric variations of a base trigger design. To select a suitable trigger design, we evaluate existing dirty-label one-to-one backdoor attacks against four critical characteristics, as detailed in Table 1. These characteristics are defined as follows: Explicit Parameters refers to whether the trigger is defined by tunable parameters (e.g., spatial coordinates, rotation angles), which is fundamental for achieving a flexible one-to-N mapping; Not Domain-Restricted indicates whether the attack is universal, meaning the trigger can be applied to any object category without relying on a specific source-to-target affinity; Stealth Preservation measures the ability to maintain imperceptibility when the parameters are adjusted across their operational range, which is crucial for avoiding detection in a multi-target setting; Robustness to Augmentation assesses the resistance to common data augmentation techniques (e.g., rotation), a practical necessity for reliable attacks when such defenses are employed.
The comparison in Table 1 highlights the shortcomings of these methods: PointBA-O is vulnerable to rotation augmentations; IRBA’s nonlinear transformations often sacrifice stealth under parameter adjustment; iBA relies on a data-driven generator lacking simple, tunable parameters; and ABA is constrained by pre-defined inter-class affinities. It is the absence of these critical flaws that makes PointBA-I the optimal choice, as it alone satisfies all four requirements. In PointBA-I, a spherical trigger fixed at a specific location can cause any point cloud containing it to be misclassified, as illustrated in Fig. 2(a). Consequently, we adopt its spherical trigger design, whose well-defined parameter space provides an ideal foundation for our parametric one-to-N attack framework.